OpenStudy (anonymous):

A bag contains 2 white marbles and 7 purple marbles. Two marbles are drawn at random. One marble is drawn and not replaced. Then a second marble is drawn. What is the probability of selecting a purple marble and then a white marble? What is the probability of selecting two white marbles? Is there a greater chance of selecting two white marbles in a row or two purple marbles in a row?

OpenStudy (anonymous):

Please help me! Thanks. -Vivid

OpenStudy (madig):

1) \[7/9*2/8= 7/36\] 2) \[2/9*1/8=1/36\] 3) 2 white marbles: \[2/9*1/8=1/36\] 2 purple marbles: \[7/9*6/8= 7/12\]